OpenStudy (anonymous):

How can i simplify this cos(-pi + x)cot(pi/2 - x)

OpenStudy (anonymous):

@dan815 Hey buddy! Wanna help? ^_^

OpenStudy (aravindg):

use property cos(-x)=cos x

OpenStudy (anonymous):

Doesnt it = -cos(x)?

OpenStudy (aravindg):

no cos is postive on I and IV quadrants ..inact cos is an even function

OpenStudy (anonymous):

Did i mistake it for sin? or is sin the same also

OpenStudy (aravindg):

sin(-u)= -sin(u) Odd cos(-u)= cos(u) Even tan(-u)= -tan(u) Odd sec(-u)= sec(u) Even csc(-u)= -csc(u) Odd cot(-u)= -cot(u) Odd where u=any real number You just have to memorize these

OpenStudy (anonymous):

Oh man buddy thanks i really needed that! ^_^

OpenStudy (aravindg):

yw :)

OpenStudy (anonymous):

so back to the question...

OpenStudy (anonymous):

i cant quite get the cot(pi/2) part down

OpenStudy (dan815):

this is what defines odd function -f(x)=f(x) like x even = f(-x)=f(x) like x^2

OpenStudy (aravindg):

it is tan x :)

OpenStudy (aravindg):

cot(pi/2 - x)=tan x

OpenStudy (anonymous):

hmmmm wait why is that?

OpenStudy (anonymous):

OH! 90degrees - x is cofunction

OpenStudy (aravindg):

just study these thoroughly http://www.mathwords.com/t/trig_identities.htm

OpenStudy (aravindg):

you can answer all these qns easily

OpenStudy (anonymous):

Okay thanks! I understand now! ^_^ I will study those

OpenStudy (aravindg):

yw :)

OpenStudy (anonymous):

Thanks @dan815 also :P

OpenStudy (dan815):